Let X and Y be jointly distributed with p.d.f:


fxy(x,y)=1/4(1+xy) ; |x|>1 |y|>1
=0. ; Otherwise
Show that X and Y are not independent but X2 and Y2 are independent
1
Expert's answer
2020-02-24T11:39:38-0500

Let X and Y be jointly distributed with p.d.f:  


"f_{X,Y}(x,y)= \begin{cases}
 {1 \over 4}(1+xy) & |x|<1, |y|<1 \\
 0 &\text{Otherwise }
\end{cases}"

Check


"\displaystyle\int_{-\infin}^{\infin}\displaystyle\int_{-\infin}^{\infin}f_{X,Y}(x,y)dydx="




"=\displaystyle\int_{-1}^{1}\displaystyle\int_{-1}^{1} {1 \over 4}(1+xy)dydx="




"=\displaystyle\int_{-1}^{1}{1 \over 4}\bigg[(y+{xy^2 \over 2})\bigg]\begin{matrix}
 1 \\
 -1 
\end{matrix}dx="

"=\displaystyle\int_{-1}^{1}{1 \over 4}(1+{x(1)^2 \over 2}-(-1+{x(-1)^2 \over 2}))dx="

"=\displaystyle\int_{-1}^{1}{1 \over 2}dx={1 \over 2}\big[x\big]\begin{matrix}
 1 \\
 -1 
\end{matrix}"

Marginal distribution of X


"g(x)=\displaystyle\int_{-\infin}^{\infin}f_{X,Y}(x,y)dy="

"=\displaystyle\int_{-1}^{1}{1 \over4}(1+xy)dy={1 \over 4}\bigg[(y+{xy^2 \over 2})\bigg]\begin{matrix}
 1 \\
 -1 
\end{matrix}="

"={1 \over 4}(1+{x(1)^2 \over 2}-(-1+{x(-1)^2 \over 2}))={1 \over 2}"

Marginal distribution of Y


"h(y)=\displaystyle\int_{-\infin}^{\infin}f_{X,Y}(x,y)dx="

"=\displaystyle\int_{-1}^{1}{1 \over4}(1+xy)dx={1 \over 4}\bigg[(x+{x^2y \over 2})\bigg]\begin{matrix}
 1 \\
 -1 
\end{matrix}="

"={1 \over 4}(1+{y(1)^2 \over 2}-(-1+{y(-1)^2 \over 2}))={1 \over 2}"

"f_{X}(x,y)= \begin{cases}
 {1 \over 2} & |x|<1, |y|<1 \\
 0 &\text{Otherwise }
\end{cases}""f_{Y}(x,y)= \begin{cases}
 {1 \over 2} & |x|<1, |y|<1 \\
 0 &\text{Otherwise }
\end{cases}"

The support set is a rectangle, so we need to check if it is true that "f_{XY}(x,y)=f_{X}(x)f_{Y}(y)," for all "(x, y)." We easily find counterexamples:


"f_{X,Y}({1 \over 2},{1 \over 2})={1 \over 4}(1+{1 \over 2}({1 \over 2}))={5 \over 16}"

"f_{X}({1 \over 2})f_{Y}({1 \over 2})={1 \over 2}({1 \over 2})={1 \over 4}"

"f_{X,Y}({1 \over 2},{1 \over 2})\not=f_{X}({1 \over 2})f_{Y}({1 \over 2})"

Therefore, X and Y are not independent.


"U=X^2, V=Y^2, 0\leq u<1, \leq v<1"

If "-1<x<0 \ and -1<y<0"

"X=-\sqrt{U}, Y=-\sqrt{V}"

The Jacobian


"J(u, v)=\begin{vmatrix}
 \partial x/\partial u & \partial x/\partial v \\
 \partial y/\partial u & \partial y/\partial v
\end{vmatrix}=""=\begin{vmatrix}
 -1/(2\sqrt{u}) & 0 \\
 0 & -1/(2\sqrt{v})
\end{vmatrix}=1/(4\sqrt{uv})"

If "0\leq x<1 \ and\ 0\leq y<1"

"X=\sqrt{U}, Y=\sqrt{V}"


"J(u, v)=\begin{vmatrix}
 \partial x/\partial u & \partial x/\partial v \\
 \partial y/\partial u & \partial y/\partial v
\end{vmatrix}=""=\begin{vmatrix}
 1/(2\sqrt{u}) & 0 \\
 0 & 1/(2\sqrt{v})
\end{vmatrix}=1/(4\sqrt{uv})"


"{1 \over 4}(1+\sqrt{uv})({1 \over 4\sqrt{uv}})= {1 \over 16}(1+{1 \over \sqrt{uv}})"

If "-1<x<0 \ and\ 0\leq y<1" or "\ 0\leq x<1 \ and\ 0\leq y<1"

"X=-\sqrt{U}, Y=\sqrt{V}"

"J(u, v)=\begin{vmatrix}
 \partial x/\partial u & \partial x/\partial v \\
 \partial y/\partial u & \partial y/\partial v
\end{vmatrix}=""=\begin{vmatrix}
 -1/(2\sqrt{u}) & 0 \\
 0 & 1/(2\sqrt{v})
\end{vmatrix}=-1/(4\sqrt{uv})"

If "0\leq x<1 \ and\ -1< y<0"

"X=\sqrt{U}, Y=-\sqrt{V}"


"J(u, v)=\begin{vmatrix}
 \partial x/\partial u & \partial x/\partial v \\
 \partial y/\partial u & \partial y/\partial v
\end{vmatrix}=""=\begin{vmatrix}
 1/(2\sqrt{u}) & 0 \\
 0 & -1/(2\sqrt{v})
\end{vmatrix}=-1/(4\sqrt{uv})"


"{1 \over 4}(1-\sqrt{uv})(-{1 \over 4\sqrt{uv}})= {1 \over 16}(1-{1 \over \sqrt{uv}})"

If "-1<x<0 \ and -1<y<0"


"f_{U,V}(u,v)= \begin{cases}
 {1 \over 16}(1+{1 \over \sqrt{uv}}) & 0<u<1,0<v<1 \\
 0 &\text{Otherwise }
\end{cases}""q(u)=\displaystyle\int_{-\infin}^{\infin}f_{U,V}(u,v)dv=""=\displaystyle\int_{0}^{1} {1 \over 16}(1+{1 \over \sqrt{uv}})dv={1 \over 16}(1+{1 \over \sqrt{u}})"

"w(v)=\displaystyle\int_{-\infin}^{\infin}f_{U,V}(u,v)du=""=\displaystyle\int_{0}^{1} {1 \over 16}(1+{1 \over \sqrt{uv}})du={1 \over 16}(1+{1 \over \sqrt{v}})"

"f_{U}(v)f_{V}(u)={1 \over 16}(1+{1 \over \sqrt{u}})({1 \over 16})(1+{1 \over \sqrt{v}})\not=""\not= {1 \over 16}(1+{1 \over \sqrt{uv}})=f_{U,V}(u,v)"

Therefore, X2 and Y2 are not independent.